find the average.
4,5,6,7,8
2,4,5,6
11,12,13,14,15,16
1
Expert's answer
2013-05-16T09:29:49-0400
The average is found as a ratio of the sum of all values divided by the number of values. So we obtain:
average(4,5,6,7,8)=(4+5+6+7+8)/5=30/5=6
average(2,4,5,6)=(2+4+5+6)/4=17/4=4.25
average(11,12,13,14,15,16)=(11+12+13+14+15+16)/6=81/6=13.5
as easy as that :)